# Transcendence and normality of complex numbers via Hurwitz continued fractions

### v2

@@ 7.2. Proof of the algorithm
  • When $\mathsf{mir}=\mathsf{mir}_{1}\mathsf{mir}_{2}$, then $a_{j_{N}}=2-i$ and $\operatorname{Re}(a_{{j_{N}}+1})\leq-1$. However, we cannot have (25) if $a_{j_{N}}=2-i$.
Figure 7. Possible forms of $\mathfrak{F}(a_{1},\ldots,a_{j_{N}+1})$ when $\mathsf{mir}$ is the identity map.
Figure 8. Possible forms of $\mathfrak{F}(a_{1},\ldots,a_{j_{N}+1})$ when $\mathsf{mir}=\mathsf{mir}_{2}$.
Case $j_{N}\geq j_{N-1}+2$. By construction, for some $m\in\mathbb{Z}$ with $|m|\geq 2$, we have
@@
and, by Remark 7.5, we may pick $\mathsf{mir}\in\mathsf{Mir}$ such that
@@
As a consequence-by the choice of $j_{N}$, Proposition 3.4 and Lemma 7.1-we have
@@
For $\mathbf{x}=(x_{n})_{n\geq 1}\in\mathscr{D}^{\mathbb{N}}$ write $\mathsf{mir}(\mathbf{x})=(\mathsf{mir}(x_{j}))_{j\geq 1}$. Let $\mathbf{d}$ be the sequence in $\mathscr{D}$ obtained from the first iteration of the algorithm applied on $\mathbf{a}^{\prime}:=(a_{n})_{n\geq j_{N-1}+2}$, then $\Lambda(\mathbf{a}^{\prime})=\overline{\Lambda}(\mathbf{d})$. If $\mathbf{b}^{\prime}:=(b_{N}(n))_{n\geq j_{N-1}+2}$ and $\mathbf{e}:=(b_{N+1}(n))_{n\geq j_{N-1}+2}$, then $\mathsf{mir}(\mathbf{a}^{\prime})=\mathbf{b}^{\prime}$, $\mathsf{mir}(\mathbf{d})=\mathbf{e}$ (see Lemma 7.2) and Lemma 7.4 yields
$\Lambda(\mathsf{mir}(\mathbf{a}^{\prime}))=\mathsf{mir}\left(\Lambda(\mathbf{a}^{\prime})\right)=\mathsf{mir}\left(\overline{\Lambda}(\mathbf{d})\right)=\overline{\Lambda}(\mathsf{mir}(\mathbf{d}))=\overline{\Lambda}(\mathbf{e})$
and (29) follows.
@@ 8. Questions for further studies
In recent years, the theory of continued fractions in certain imaginary quadratic number fields has attracted some attention [16, 17, 18]. Although our work uses the symmetries of the HCF process, we believe that a general procedure can be obtained to cover the continued fraction algorithms used in the papers mentioned above. Hence, one of the open questions is to explore:
@@
Thermodynamic formalism studies measures of maximal entropy and equilibrium measures of dynamical systems. This area has interesting interpretations in physics and (fractal) dimension theory (like Bowen’s formula [7]). Thermodynamic formalism for non-finite alphabets is an active area with interesting connections to surface dynamics (see the survey [33]). Unlike the finite-alphabet theory, techniques and results are dependent on the class of systems.
@@
Acknowledgements. The authors thank anonymous referees for the careful reading of the manuscript and a number of comments that have improved the quality of this work. We also thank Nikita Shulga for his comments and suggestions

### v3

@@ 7.1. Auxiliary lemmas
  4. iv.
    When $\mathsf{mir}=\mathsf{mir}_{1}\mathsf{mir}_{2}$, then $a_{j_{N}}=2-i$ and $\operatorname{Re}(a_{{j_{N}}+1})\leq-1$. However, we cannot have (31) for $k=N$ if $a_{j_{N}}=2-i$.
Figure 7. Possible forms of $\mathfrak{F}(a_{1},\ldots,a_{j_{N}})$ when $\mathsf{mir}$ is the identity map.
Figure 8. Possible forms of $\mathfrak{F}(a_{1},\ldots,a_{j_{N}})$ when $\mathsf{mir}=\mathsf{mir}_{2}$.
Case $j_{N}\geq j_{N-1}+2$. By the definition of $\mathbf{b}_{N}$, for some $m\in\mathbb{Z}$ with $|m|\geq 2$, we have
@@
By Remark 7.4, there is some $\mathsf{mir}\in\mathsf{Mir}$ such that
@@
By the choice of $j_{N}$, Proposition 3.4 and Lemma 7.1, we conclude
@@
For $\mathbf{x}=(x_{n})_{n\geq 1}\in\mathscr{D}^{\mathbb{N}}$, put $\mathsf{mir}(\mathbf{x})=(\mathsf{mir}(x_{j}))_{j\geq 1}$. Let $\mathbf{d}=(d_{j})_{j\geq 1}$ be obtained from the first iteration of Procedure 4.3 applied on $\mathbf{a}^{\prime}:=(a_{n})_{n\geq j_{N-1}+2}$, then $\Lambda(\mathbf{a}^{\prime})=[0;d_{1},d_{2},\ldots]$. If $\mathbf{b}^{\prime}:=(b_{N}(n))_{n\geq j_{N-1}+2}$ and $\mathbf{e}:=(b_{N+1}(n))_{n\geq j_{N-1}+2}$, then $\mathsf{mir}(\mathbf{a}^{\prime})=\mathbf{b}^{\prime}$, $\mathsf{mir}(\mathbf{d})=\mathbf{e}$ (see Lemma 7.2) and direct computations give
$\overline{\Lambda}(\mathsf{mir}(\mathbf{a}^{\prime}))$ $=\mathsf{mir}\left(\overline{\Lambda}(\mathbf{a}^{\prime})\right)$
$=\mathsf{mir}\left([0;d_{1},d_{2},\ldots]\right)=[0;\mathsf{mir}(d_{1}),\mathsf{mir}(d_{2}),\ldots]=[0;e_{1},e_{2},\ldots].$
Equation (36) follows. We can, thus, conclude Assertions A and B.
Conclusion of the proof.
Lastly, we show Assertion C. Let $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 0}$ be the sequences defined as in (6) corresponding to $\mathbf{b}$. Then, $\frac{p_{n}}{q_{n}}\to\overline{\Lambda}(\mathbf{b})$ as $n\to\infty$ and $(|q_{n}|)_{n\geq 0}$ is strictly increasing. Take $N\in\mathbb{N}$. By Remark 7.4, there is some $\mathsf{mir}^{\prime}\in\mathsf{Mir}$ such that
$\xi_{N+1}:=[0;b_{N+1}(j_{N}+2),b_{N+1}(j_{N}+3),\ldots]=\mathsf{mir}^{\prime}\left([0;a_{j_{N}+2},a_{j_{N}+3},\ldots]_{\mathbb{C}}\right)\in\overline{\mathfrak{F}}.$
Since $(b_{1},\ldots,b_{j_{N}+1})=(b_{N+1}(1),\ldots,b_{N+1}(j_{N}+1))$, the above argument and the elementary theory of continued fractions [27, Theorems 2 and 5] yield
$\Lambda(\mathbf{a})=[0;b_{N+1}(1),b_{N+1}(2),\ldots]=\frac{p_{j_{N}+1}}{q_{j_{N}+1}}+\frac{(-1)^{j_{N}+1}\xi_{N+1}}{q_{j_{N}+1}\left(1+\frac{q_{j_{N}}}{q_{j_{N}+1}}\xi_{N+1}\right)}.$
Then, using $|\xi_{N+1}|\leq\sqrt{2}$, we have
$\left|\Lambda(\mathbf{a})-\frac{p_{j_{N}+1}}{q_{j_{N}+1}}\right|\leq\frac{2}{2-\sqrt{2}}\,\frac{1}{|q_{j_{N}+1}|^{2}}\to 0\;\text{ as }\;N\to\infty,$
and $\overline{\Lambda}(\mathbf{b})=\Lambda(\mathbf{a})$ follows. With the three assertions proven, Proposition 4.1 is established.
@@ 8. Questions for further studies
In recent years, the theory of continued fractions in certain imaginary quadratic number fields has attracted some attention [16, 17, 18]. Although our work relies on the symmetries of the HCF process, we believe that a general procedure can be obtained to cover the continued fraction algorithms used in the papers mentioned above. Hence, one of the open questions is to explore:
@@
Thermodynamic formalism studies measures of maximal entropy and equilibrium measures of dynamical systems. This area has interesting interpretations in physics and (fractal) dimension theory (like Bowen’s formula [7]). Thermodynamic formalism for non-finite alphabets is an active area with interesting connections to surface dynamics (see the survey [33]). Unlike the finite-alphabet theory, techniques and results depend on the class of systems.
@@
Acknowledgements. The authors thank anonymous referees for the careful reading of the manuscript and several comments that have improved the quality of this work. We also thank Nikita Shulga and Ben Ward for their comments and suggestions.